OpenStudy (anonymous):

A heated piece of metal cools according to the function c(x) = (.5)x − 10, where x is measured in hours. A device is added that aids in cooling according to the function h(x) = −x − 1. What will be the temperature of the metal after five hours? −6° Celsius 26° Celsius 32° Celsius 38° Celsius

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@shadowfiend @Preetha

OpenStudy (solomonzelman):

I think all you have to do is to plug -5 for x into h(x) = −x − 1 h(x)=-5-1 h(x)=-6 but I'm not sure about this.

OpenStudy (anonymous):

That's what I thought, so I'll put it and hope we're right!

OpenStudy (solomonzelman):

Tell me when you submit it.

OpenStudy (anonymous):

I have 6 more questions I'm trying to figure out before I can submit, do you think you can help?

OpenStudy (solomonzelman):

I can try.

OpenStudy (anonymous):

This is the first one.

OpenStudy (solomonzelman):

I would hit both sides with a log.\[\log(1/(2\sqrt{2})=(m+2)\log4\]

OpenStudy (solomonzelman):

I need to think....

OpenStudy (solomonzelman):